HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 99     Entry time: Wed Apr 2 23:40:23 2025
Author: JEL, ZN 
Category: General 
Subject: Screenshots taken just as beam came back 
We are seeing spills being received but no peaks are seen in the analyzers. 

We are seeing several intermittent losses of beam. Awaiting action to see if it stabilizes.


23.51 Beam is back and we see the nice spectrum in the analyzer.
Attachment 1: 2025-04-02_23-39-29-655.png  627 kB  Uploaded Thu Apr 3 00:40:47 2025  | Hide | Hide all
2025-04-02_23-39-29-655.png
Attachment 2: 2025-04-02_23-37-10-651.png  245 kB  Uploaded Thu Apr 3 00:40:51 2025  | Hide | Hide all
2025-04-02_23-37-10-651.png
Attachment 3: 2025-04-02_23-36-58-664.png  429 kB  Uploaded Thu Apr 3 00:40:57 2025  | Hide | Hide all
2025-04-02_23-36-58-664.png
ELOG V3.1.5-fc6679b